OpenStudy (anonymous):

help on finding inverse functions! f(x) = x^3 - 8

OpenStudy (irishboy123):

write it as \(y = x^3 - 8\) ad then find \(x = \ ...\) in terms of y

OpenStudy (anonymous):

i know how to do that, i added 8 to both sides but im stuck on what to do with the cubed x

OpenStudy (tkhunny):

Have you considered a Cube Root?

OpenStudy (anonymous):

whats that

OpenStudy (irishboy123):

\(\sqrt[3]{27} = 3\) because \(3 \times 3 \times 3 = 27\)

OpenStudy (irishboy123):

\(\sqrt[3]{whatever}\) is the cube root of whatever

OpenStudy (anonymous):

oh, so it would be f^-1 (x) = 3√x + 8 with the radical going over both x and 8 or just x?

OpenStudy (tkhunny):

Pretty much. Or, you can use a 1/3 exponent.

OpenStudy (anonymous):

its a one-to one function right

OpenStudy (tkhunny):

In this case, it is.

OpenStudy (anonymous):

im confused because i know its one of the 2 choices, both have f(-1)x = 3√x + 8, but one has the radicand or whatever its called over the x ONLY, and the other has the radicand over the whole expression (x+8) which one is it?

OpenStudy (tkhunny):

You should not be confused. Many exams have fake answers that might be plausible on first inspection. You must pick the right one.

OpenStudy (anonymous):

its the one i picked

OpenStudy (anonymous):

with the radicand going over the whole thing

OpenStudy (tkhunny):

If you took the cube root of the whole thing, then yes. Did you?

OpenStudy (anonymous):

no only for x^3

OpenStudy (tkhunny):

\(y = x^{3} - 8\) Swap \(x = y^{3} - 8\) Solve \(x + 8 = y^{3}\) \(\sqrt[3]{x + 8} = y\) Watch yourself work. Don't guess.